### v1

# On energy and its positivity in spacetimes with an expanding flat de Sitter background

Abstract.
The positive energy theorems are a fundamental pillar in mathematical general relativity. Originally proved by Schoen–Yau and later Witten, these theorems were established for asymptotically flat manifolds where the metric tends to the standard Euclidean metric and whose second fundamental form decays to zero at infinity. This ansatz on the metric and second fundamental form is motivated by the desire to model an isolated gravitational system with a Minkowski space background for the spacetime. However, actual astrophysical massive objects are not truly isolated but rather exist within an expanding cosmological universe, where the second fundamental form is umbilic. With this in mind, we seek a notion of energy for initial data sets with an umbilic second fundamental form. In this work, we present a definition of energy in such an expanding cosmological setting. Instead of Minkowski space, we take de Sitter space as the background spacetime, which, when written in flat-expanding coordinates, is foliated by umbilic hypersurfaces each isometric to Euclidean 3-space. This cosmological setting necessitates a quasi-local energy definition, as the presence of a cosmological horizon in de Sitter space obstructs a global one. We define energy in this quasi-local setting by adapting the Liu–Yau energy to our framework and establish positivity of this energy for certain bounded values of the cosmological constant.
1. Introduction
Let $(\mathscr{M},\mathscr{g})$ be a four-dimensional spacetime, i.e., a time-oriented Lorentzian manifold. An initial data set is a quadruple $(M,g,k,\lambda)$, where $M$ is a spacelike hypersurface, $g$ is the metric induced on $M$, $k$ is the second fundamental form of $M$ in $\mathscr{M}$, and $\lambda$ is a constant. We note that the pair $(g,k)$ may depend on $\lambda$. In what follows we assume that $\lambda\geq 0.$ When $\lambda=0$, we adopt the classical notation $(M,g,k)$. In particular, if $u$ is the future-directed unit timelike vector field, then, in our convention,
$k(X,Y)=\langle\nabla_{X}u,Y\rangle,$ (1)
where $X,Y$ are tangent vector fields on $M$ and $\nabla$ is the spacetime Levi-Civita connection. With this convention, a positive mean curvature of the slice corresponds to expansion towards the future.
The initial data set must satisfy the Einstein constraint equations, which arise from the Gauss–Codazzi decomposition of the Einstein field equations. In the presence of a cosmological constant $\Lambda=3\lambda^{2}$, these constraints take the form:
$2\rho+2\Lambda$ $=R_{g}+({\rm tr}_{g}k)^{2}-|k|_{g}^{2},$ (2)
$J$ $={\rm div}_{g}\left[k-({\rm tr}_{g}k)g\right],$
where $R_{g}$ is the scalar curvature of the induced metric $g$, $\rho$ denotes the energy density, and $J$ is the momentum density of the matter fields.
An initial data set is said to satisfy the dominant energy condition with respect to $\Lambda$ if the inequality
$\rho\geq|J|_{g}$ (3)
holds.
We remark that the dominant energy condition used here is derived from the spacetime dominant energy condition applied to the stress-energy tensor $\mathscr{T}=\mathcal{G}+\Lambda\mathscr{g},$ where $\mathcal{G}$ is the Einstein tensor associated with $\mathscr{g}$. Specifically, we require that for every future-directed timelike vector $\mathscr{v}$, the vector $\mathscr{T}(\mathscr{v},\cdot)^{\sharp}$ is also future-directed timelike. Consequently, this dominant energy condition depends explicitly on $\Lambda$. This condition ensures that energy density dominates momentum flux, reflecting the physical requirement that energy flows cannot exceed the speed of light.
If $\Lambda\geq 0$ and the dominant energy condition with $\Lambda$ holds, then the geometrical energy density $\mu$ satisfies
$\mu:=\rho+\Lambda\geq|J|_{g}.$ (4)
Note that this is the dominant energy condition without $\Lambda$. Moreover, $\mu$ is determined entirely by the geometry of the initial data:
$\mu=\frac{1}{2}\bigg(R_{g}+({\rm tr}_{g}k)^{2}-|k|_{g}^{2}\bigg).$ (5)
Noether’s theorem (see [Arn89]), loosely speaking, states that conserved quantities correspond to symmetries of the system. Therefore, to define conserved quantities such as energy in general relativity, we look for global symmetries of the spacetime. These symmetries are generated by Killing vector fields, which are vector fields that preserve the metric along their flow (i.e. they generate isometries of the spacetime). In particular, a timelike Killing vector field corresponds to a time-translation symmetry and, by Noether’s theorem, gives rise to a conserved energy. However, most spacetimes do not possess such global symmetries. This is why asymptotically Minkowskian spacetimes are of particular interest: in these spacetimes the geometry approaches Minkowski space at infinity, where a timelike Killing vector field exists. This asymptotic symmetry provides the structure needed to define a conserved energy at infinity.
The ADM energy [ADM61] $E_{\mathrm{ADM}}$ plays a central role in the analysis of asymptotically flat spacetimes. It is derived from the Hamiltonian formulation of general relativity under the assumption of a fixed Minkowski background. It provides a notion of total energy for an isolated gravitational system and is invariant under asymptotically flat coordinate transformations [Bar86, Chr88]. This invariance guarantees that the energy is well-defined independently of the observer’s coordinate choice at infinity.
The ADM energy is defined for initial data sets $(M,g,k,\lambda=0)$ that satisfy specific fall-off conditions at spatial infinity, namely:
$g_{ij}=\delta_{ij}+O_{2}(|x|^{-\tau}),\qquad k_{ij}=O_{1}(|x|^{-1-\tau}),$ (6)
where $\tau>\frac{1}{2}$. These asymptotic conditions ensure that the metric $g$ approaches the flat Euclidean metric and that the extrinsic curvature $k$ decays appropriately. Assuming integrability of $\mu$ and $J$, the energy is then defined as
$E_{\mathrm{ADM}}=\frac{1}{16\pi}\lim_{r\to\infty}\int_{S_{r}}\sum_{i,j}\left(\partial_{i}g_{ij}-\partial_{j}g_{ii}\right)\frac{x_{j}}{r}\,d\mu_{\delta},$ (7)
where $d\mu_{\delta}$ denotes the area measure on the coordinate sphere $S_{r}=\{|\mathbf{x}|=r\}$ induced by the Euclidean metric $\delta$. We remark that in this setting also the linear momentum and ADM mass would be well defined.
A fundamental result in mathematical relativity is the Positive Energy Theorem. It was first proved for the time symmetric case $k\equiv 0$ independently by Schoen and Yau [SY79] and Witten [Wit81]. It was later extended for the case of a general $k$ with appropriate asymptotic decay by Schoen and Yau [SY81a, SY81b]. The theorem states that, under the assumption that the dominant energy condition holds, the ADM energy is non-negative:
$E_{\mathrm{ADM}}\geq 0.$ (8)
Equality occurs if and only if the initial data $(M,g,k)$ arises from a spacelike slice of flat Minkowski spacetime.
Extensions of the positive energy theorem exist for other asymptotic geometries. In the case $\Lambda<0$, corresponding to asymptotically hyperbolic manifolds, the notion of energy was introduced by Wang [Wan01] and further developed by Chruściel and Herzlich [CH03]. The Riemannian positive mass theorem in this setting was proven under the assumption that the manifold is spin in [Wan01, CH03]. The positivity of the energy in the hyperboloidal, non-spin case has been established by Sakovich (see [Sak21] and references therein).
When $\Lambda>0$, the situation becomes more delicate. A Positive Energy Theorem in this setting has been established by Borghini and Mazzieri [BM18, BM20] for compact static metrics. We also note the connection with rigidity and the Min–Oo conjecture (see [BMN11], where it was shown to be false). In contrast with [BM18, BM20], our focus is on umbilic initial data sets, which are not static.
The question we want to address is how to describe isolated gravitational systems within an expanding universe. Observational evidence of the accelerated expansion of the universe strongly supports a cosmological model with a positive cosmological constant $\Lambda>0$.
To study isolated systems in this framework, we adopt a background spacetime model that reflects this expansion. The flat expanding de Sitter model provides a natural choice, as it captures the large-scale behavior of a universe dominated by $\Lambda$-vacuum energy while retaining the mathematical simplicity of spatial flatness. Our goal is to find a definition of energy in this setting and establish a corresponding positivity result.
[…]
Figure 1. One-half of de Sitter space is covered by the spatially flat FLRW model with an exponentially growing scale factor. In flat-expanding coordinates, each $t$-slice is isometric to the Euclidean space $\mathbb{R}^{3}$ and $\partial_{t}$ is Killing but timelike only for values within the cosmological horizon, i.e. for $r<1/\lambda$. Compare with [HE73, page 127]. For an in-depth analysis of how the flat-expanding coordinates relate to the global coordinates of de Sitter space, see [GLQ23].
Recall that the De Sitter spacetime is a maximally symmetric solution to Einstein’s equations with a positive cosmological constant $\Lambda>0$. It serves as a model for a universe undergoing accelerated expansion, which is supported by current observations. In comoving coordinates, the metric for the de Sitter spacetime takes the form
$\mathscr{M}_{0}=\mathbb{R}\times\mathbb{R}^{3}\quad\text{ and }\quad\mathscr{g}_{0}=-d\tau^{2}+e^{2\lambda\tau}(d\rho^{2}+\rho^{2}d\Omega^{2}),$ (9)
where
$\lambda:=\sqrt{\frac{\Lambda}{3}}.$ (10)
Equation (9) is the standard spatially flat FLRW model with scale factor $a(\tau)=e^{\lambda\tau}$.
In flat-expanding coordinates, $t=\tau$, $r=e^{\lambda\tau}\rho$, the metric takes the form:
$\mathscr{g}_{0}=-\left(1-r^{2}\lambda^{2}\right)dt^{2}-2\lambda r\,drdt+dr^{2}+r^{2}d\Omega^{2}.$ (11)
Constant $t$ slices of the above metric are spatially flat and umbilic, with second fundamental form
$k_{0}=\lambda\delta.$ (12)
Therefore, this particular foliation of de Sitter spacetime yields an ideal background to model spatially flat systems which undergo cosmological expansion.
In flat-expanding coordinates, the vector field $\partial_{t}$ is a Killing vector field; however, it ceases to be timelike for $r$-values greater than the cosmological horizon, $r=1/\lambda$. This leads to complications with an asymptotic definition of energy. The presence of this cosmological horizon suggests the need for a shift from global to quasi-local definitions of energy.
Quasi-local notions of energy play a central role in mathematical relativity, as they provide a framework for defining and analyzing the energy content of finite spacetime regions (see [Lee19] for a survey of various notions of quasi-local energy and their properties).
In the quasi-local setting, one considers an initial data set $(\Omega,g,k,\lambda)$, where $\Omega$ is compact with some boundary component [^1] In general, $\Omega$ can have multiple boundary components. The perspective here is that $\Sigma$ is an outer boundary component that encompasses multiple inner boundary components that could represent black hole horizons. $\Sigma:=\partial\Omega$ of positive Gauss curvature. By the Weyl embedding theorem [Nir53, Pog52], there is an isometric embedding of $\Sigma$ within Euclidean space $\mathbb{R}^{3}$ which bounds a compact and convex subset $\Omega_{0}$. Let $H_{0}$ and $H$ denote the mean curvatures of $\Sigma$ within $\mathbb{R}^{3}$ and $\Omega$, respectively. (In our convention, the mean curvature is the trace of the second fundamental form with respect to the outward unit normal.) Then we define the energy $E_{\lambda}$ as a direct generalization of the Liu–Yau energy [LY03, LY06]:
$E_{\lambda}:=\frac{1}{8\pi}\int_{\Sigma}\left(\sqrt{H_{0}^{2}-4\lambda^{2}}\,-\,\sqrt{H^{2}-({\rm tr}_{\Sigma}k)^{2}}\right)d\mu.$ (13)
As for the Liu–Yau energy, $E_{\lambda}$ can be expressed as the difference between the norms of the codimension-two mean curvature vectors in the relevant spacetimes. Indeed it’s not hard to see that
$E_{\lambda}=\frac{1}{8\pi}\int_{\Sigma}\big(|\vec{H}_{0}|\,-\,|\vec{H}|\big)\,d\mu.$ (14)
Here $\vec{H}$ is the mean curvature vector of $\Sigma$ (assumed to be spacelike) within the original spacetime $(\mathscr{M},\mathscr{g})$ containing $(\Omega,g,k,\lambda)$, while $\vec{H}_{0}$ is the mean curvature vector of $\Sigma$ within the de Sitter spacetime $(\mathscr{M}_{0},\mathscr{g}_{0})$ with $\Lambda=3\lambda^{2}$ containing the compact initial data set $(\Omega_{0},\delta,k_{0}=\lambda\delta,\lambda)$ within a constant $t$ slice. Both mean curvature vectors are spacelike in this setting. See figure 2 for a comparison between the Liu-Yau energy and $E_{\lambda}$.
The goal of this paper is to establish positivity results for $E_{\lambda}$. Analogous to the classical Liu–Yau framework, this requires ensuring that the data under consideration can be embedded isometrically into $\mathbb{R}^{3}$. Moreover, for the definition to be physically meaningful in a de Sitter background, the embedding must lie within the cosmological horizon $r<1/\lambda$ where the Killing field $\partial_{t}$ is timelike (see figure 2).
In Section 2 we first state a general conjecture concerning the positivity of $E_{\lambda}$. The conjecture naturally leads to a two-fold problem:
  1. (1)
    a purely geometric question, concerning the conditions under which a given surface $\Sigma$ can be isometrically embedded inside the cosmological horizon;
  2. (2)
    a question regarding the positivity of the energy formula given the geometric data.
In this work, we address both issues and provide partial answers for small values of $\lambda.$
We remark that the observed value for the cosmological constant is, in fact, small. Using the observations made by [Pla20], we have $\Lambda_{\rm obs}=1.10\times 10^{-52}$ ${\rm m}^{-2}$. For this value, we have $\lambda_{\rm obs}=6.05\times 10^{-27}$ ${\rm m}^{-1}$. The cosmological horizon at this observed value has a radius of $r=1/\lambda_{\rm obs}=1.43\times 10^{26}$ $\rm m$}, which is about 100,000 times greater than the diameter of the Milky Way.
The structure of the paper is as follows. In Section 2 we state the main conjecture and our main results, Theorem 2.3 and Corollary 2.4. Section 3 reviews the relevant notions of quasi-local energy developed in a Minkowskian background, together with the associated positivity results. In Section 4 we provide the relevant lemmas and prove our main results. Section 5 explores examples of our results. Section 5.1 reviews the Schwarzschild–de Sitter spacetime in flat expanding coordinates and computes the quantities appearing in Theorem 2.3. Section 5.2 constructs many umbilic initial data sets $(\Omega,g,k=\lambda g,\lambda)$ satisfying the $\Lambda$-vacuum constraint equations (2) where our results apply. Finally, in Section 6 we summarize our findings and discuss future directions.
$\Sigma$$\mathcal{I}^{+}\mathcal{I}^{-}$$\Sigma$$\mathcal{I}^{+}\mathcal{I}^{-}$
Figure 2. Left: the Penrose diagram of Minkowski spacetime. To compute the Liu–Yau energy, the surface $\Sigma$ is isometrically embedded in a constant $t$-slice of Minkowski spacetime. Right: the Penrose diagram of de Sitter spacetime. The surface is also isometrically embedded in a constant $t$-slice of the flat de Sitter spacetime and lies within the cosmological horizon (shown in red).
2. Main conjecture and theorem
As explained above, the presence of a cosmological horizon in the flat-expanding coordinates of de Sitter space suggests the need to replace the usual notion of total energy at infinity with a quasi-local one. Motivated by this, we introduce the energy $E_{\lambda}$ defined in (13), which is a direct generalization of the Liu–Yau quasi-local energy [LY03, LY06] to the flat-expanding patch of de Sitter space. The relevant definitions and properties of the classical Liu–Yau construction will be recalled in Section 3.
Our goal is to establish positivity results for $E_{\lambda}$. In analogy with the classical Liu–Yau framework, this requires verifying that the data under consideration can be embedded isometrically into $\mathbb{R}^{3}$. Furthermore, for the definition to be physically meaningful in the de Sitter case, the embedding must lie inside the cosmological horizon of the background spacetime (see Figure 2).
Consider an initial data set $(\Omega,g,k,\lambda)$ where $\Omega$ is compact with some boundary component $\Sigma=\partial\Omega$ and $\lambda=\sqrt{\Lambda/3}\geq 0$, where $\Lambda$ is the fixed cosmological constant of the background de Sitter space. Assume that $\Sigma$ is a closed, 2-dimensional surface of positive Gauss curvature. By the Weyl embedding theorem, which was solved independently by Nirenberg and Pogorelov [Nir53, Pog52], $\Sigma$ can be isometrically embedded into the Euclidean space $\mathbb{R}^{3}$ with the image of $\Sigma$ as the boundary of a compact and convex domain $\Omega_{0}\subset\mathbb{R}^{3}$; moreover, this embedding is unique up to Euclidean rigid motions. Let $H_{0}$ be the mean curvature of the embedding in $\mathbb{R}^{3}$ with respect to the outward normal. In the initial data $\Omega$, let $H$ be the mean curvature of $\Sigma$ with respect to outward normal.
Our first conjecture states that the energy $E_{\lambda}$ should be nonnegative whenever the initial data set satisfies the dominant energy condition with respect to $\Lambda$ and the Weyl embedding of $\Sigma$ in $\mathbb{R}^{3}$ lies within the cosmological horizon; as in Liu-Yau [LY03, LY06], we also require that the mean curvature vectors of $\Sigma$ are spacelike within $\Omega$ and the Weyl embedding in $\mathbb{R}^{3}$ so that $E_{\lambda}$ is well defined.
Conjecture 2.1.
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
  1. (1)
    the Weyl embedding of $\Sigma$ in $\mathbb{R}^{3}$ lies entirely inside the cosmological horizon, $r<1/\lambda$;
  2. (2)
    $H_{0}^{2}-4\lambda^{2}>0$;
  3. (3)
    $H^{2}-({\rm tr}_{\Sigma}k)^{2}>0$;
  4. (4)
    the initial data set $(\Omega,g,k,\lambda)$ satisfies the dominant energy condition with $\Lambda=3\lambda^{2}$.
Then the quasi-local energy
$E_{\lambda}:=\frac{1}{8\pi}\int_{\Sigma}\left(\sqrt{H_{0}^{2}-4\lambda^{2}}\,-\,\sqrt{H^{2}-({\rm tr}_{\Sigma}k)^{2}}\right)d\mu$ (15)
is non-negative:
$E_{\lambda}\geq 0.$ (16)
In the perfectly umbilic setting, $k=\lambda g$, the question becomes purely Riemannian. Indeed, from the first constraint equation, the dominant energy condition with $\Lambda=3\lambda^{2}$ is equivalent to requiring nonnegative of scalar curvature, $R_{g}\geq 0$. This leads to the purely Riemannian version of Conjecture 2.1.
Conjecture 2.2.
Let $(\Omega,g,k=\lambda g,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Assume that:
  1. (1)
    the Weyl embedding of $\Sigma$ in $\mathbb{R}^{3}$ lies entirely inside the cosmological horizon, $r<1/\lambda$;
  2. (2)
    $H_{0}^{2}-4\lambda^{2}>0$;
  3. (3)
    $H^{2}-4\lambda^{2}>0$;
  4. (4)
    $R_{g}\geq 0$.
Then the quasi-local energy
$E_{\lambda}=\frac{1}{8\pi}\int_{\Sigma}\left(\sqrt{H_{0}^{2}-4\lambda^{2}}\,-\,\sqrt{H^{2}-4\lambda^{2}}\right)d\mu$ (17)
is non-negative:
$E_{\lambda}\geq 0.$ (18)
Addressing these conjectures involve two challenges. The first is geometric in nature: one would like to determine under which conditions the surface $\Sigma$ can be isometrically embedded inside the cosmological horizon. The second concerns the positivity of the energy $E_{\lambda}$ itself. We will show that both of these concerns can be resolved for some values of $\lambda$.
We start by introducing a constant $\alpha_{\lambda}=\alpha(g,k,\lambda)$ as:
$\alpha_{\lambda}=\begin{cases}\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}&\text{if }\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\\ \min\left\{\sqrt{\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right)},\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}\right\}&\text{if }\sqrt{K_{\min}}\geq\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\end{cases}.$ (19)
In this expression, $K_{\min}$ is the minimum of the Gauss curvature $K$ on $\Sigma$, $r_{\Sigma}$ is the area radius of $\Sigma$, and $E_{\rm LY}$ is the Liu-Yau energy of the initial data (see equation (24) in the next section). We note that the definition of $\alpha_{\lambda}$ does not involve $\lambda$ explicitly; however, an implicit dependence may occur if either the metric $g$ or the second fundamental form $k$ (and hence the Liu–Yau energy $E_{\rm LY}$) depends on $\lambda$.
Our main result, Theorem 2.3, addresses Conjecture 2.1 for some values of $\lambda$. Then we will find that $E_{\lambda}\geq 0$ so long as $\lambda\leq\alpha_{\lambda}$. Recognize that $\alpha_{\lambda}$ is nonnegative whenever the initial data set satisfies the dominant energy condition with respect to $\Lambda\geq 0$. Moreover $\alpha_{\lambda}$ is strictly positive if the geometrical energy density $\mu$ is strictly positive, which follows from the rigidity statement of Theorem 3.2.
Theorem 2.3.
Let $(\Omega,g,k,\lambda)$ for $\lambda\geq 0$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}$ holds. If $\lambda\leq\alpha_{\lambda}$, where $\alpha_{\lambda}$ is given by (19), then the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and the energy (15) is nonnegative:
$E_{\lambda}\geq 0.$ (20)
Moreover, if $0<\lambda<\alpha_{\lambda}$, then $E_{\lambda}>0$; if $\lambda=\alpha_{\lambda}>0$ and $E_{\lambda}=0$, then $\Sigma$ is isometric to a round sphere.
Remark.
The condition $\lambda\leq\alpha_{\lambda}$ may be always satisfied in certain settings; an explicit example is provided by the Schwarzschild–de Sitter case discussed in Section 5.1.
Remark.
We note that Theorem 2.3 holds for any choice of the second fundamental form $k$. In particular, we can adopt two different points of view:
  1. (1)
    Fixed geometric data. One may fix an initial data set $(\Omega,g,k)$ that is completely independent of $\lambda$. In this case, $\lambda$ acts as an external parameter, and $\alpha_{\lambda}$ depends only on the intrinsic geometry of $\Sigma$ and on the Liu–Yau energy $E_{\rm LY}$ computed from the fixed data and hence is independent of $\lambda$. Therefore the theorem provides an interval $0,\lambda_{\max}]$, where $\lambda_{\max}=\alpha$, of admissible values of $\lambda$ for which the data embeds within the cosmological horizon and the quasi-local energy $E_{\lambda}$ is nonnegative. In this case, the rigidity part of Theorem 2.3 can be understood as if $\lambda=\lambda_{\rm max}>0$ and $E_{\lambda}=0$, then $\Sigma$ is isometric to a round sphere.
  2. (2)
    $\lambda$-dependent data. Alternatively, one may consider a $\lambda$-dependent family of data $(\Omega,g(\lambda),k(\lambda))$, for instance determined by the constraint equations or by prescribing an umbilic second fundamental form. In this case, both $E_{\rm LY}$ and $\alpha_{\lambda}$ vary with $\lambda$, and the theorem can be applied pointwise by verifying the condition $\lambda\leq\alpha_{\lambda}$ for each value of $\lambda$.
We now address Conjecture 2.2, the Riemannian version of Conjecture 2.1. In particular, we show that there is a given interval of values of $\lambda>0$ for which the data correctly embeds within the cosmological horizon and has nonnegative energy $E_{\lambda}$. This interval will be only dependent on the geometric data $(\Omega,g)$.
Corollary 2.4.
Let $(\Omega,g)$ be a compact Riemannian manifold of nonnegative scalar curvature with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Then there exists a constant $\lambda_{\max}\geq 0$ only dependent on $(\Omega,g)$ such that, for any $\lambda\in[0,\lambda_{\rm max}]$, the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and if the mean curvature of $\Sigma$ in $(\Omega,g)$ satisfies $H>2\lambda$, then the energy (17) of the initial data set $(\Omega,g,k=\lambda g,\lambda)$ is well defined and nonnegative:
$E_{\lambda}\geq 0.$ (21)
Moreover, it’s strictly positive if $0<\lambda\leq\lambda_{\rm max}$.
Remark.
The constant $\lambda_{\max}$ in Corollary 2.4 is defined as $\alpha_{\lambda}$ in (19) by substituting $E_{\rm LY}$ with its Brown–York counterpart $E_{\rm BY}$ everywhere. Consequently, by the rigidity statement of Shi–Tam [ST02] (see Theorem 3.1 below), $\lambda_{\rm max}>0$ whenever $(\Omega,g)$ is not a subset of Euclidean space $\mathbb{R}^{3}$. For specific examples of this corollary, see Section 5.2.
3. Positivity Results for Quasi-Local Energies with a Minkowski background
In this section we provide some motivation for the quasi-local energy $E_{\lambda}$ defined by (13). Briefly, the expression for $E_{\lambda}$ is a generalization of the Brown–York and Liu–Yau definitions of energy but adapted to the setting of an expanding flat de Sitter background instead of a Minkowski background.
We begin by recalling the classical result concerning isometric embeddings of positively curved 2-surfaces. Let $\Sigma$ be a closed surface with a Riemannian metric of positive Gauss curvature. Then there exists an isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$, known as the Weyl embedding, which bounds a compact convex region $\Omega\subset\mathbb{R}^{3}$. Moreover, the embedding is unique up to Euclidean rigid motions [Nir53, Pog52]. In particular, the mean curvature of the isometric embedding is uniquely determined by the intrinsic metric. Using this fact, Brown and York proposed the following quasi-local mass:
Definition 3.1 (Brown–York Energy).
Let $\Omega$ be a compact Riemannian 3-manifold with nonempty boundary $\Sigma=\partial\Omega$ of positive Gauss curvature. Let $H$ be the mean curvature of $\Sigma$ in $\Omega$ with respect to the outward normal, and let $H_{0}$ the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Brown–York energy of $\Sigma$ is defined as
$E_{\rm BY}:=\frac{1}{8\pi}\int_{\Sigma}\left(H_{0}\,-H\right)d\mu.$ (22)
The positivity of the Brown–York energy was established by Shi and Tam [ST02] assuming the dominant energy condition and mean convexity.
Theorem 3.1 ([ST02]).
Suppose $\Omega$ is a compact Riemannian 3-manifold with non-negative scalar curvature and nonempty boundary $\Sigma=\partial\Omega$ with positive Gauss curvature and mean curvature. Then the Brown–York energy $E_{\rm BY}$ of $\Sigma$ is non-negative and equals zero if and only if $\Omega$ is isometric to a domain in the Euclidean space $\mathbb{R}^{3}$ and $\Sigma$ has only one component.
Liu and Yau [LY03, LY06] generalized this to a Lorentzian setting, incorporating the full spacetime mean curvature vector.
Definition 3.2 (Liu–Yau Energy).
Let $\Sigma$ be an closed spacelike 2-surface in a spacetime $\mathscr{M}$, and suppose $\Sigma$ has positive Gauss curvature. Let $|\vec{H}|$ denote the norm of the spacetime mean curvature vector of $\Sigma$, and let $H_{0}$ be the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Liu–Yau energy of $\Sigma$ is defined as
$E_{\mathrm{LY}}:=\frac{1}{8\pi}\int_{\Sigma}\big(H_{0}\,-\,|\vec{H}|\big)\,d\mu.$ (23)
If we assume that the surface $\Sigma$ is the boundary of a compact initial data set $(\Omega,g,k)$, then the length of the spacetime mean curvature vector $\vec{H}$ can be computed in terms of the initial data set. Indeed, if the mean curvature vector is spacelike, then the Liu–Yau energy can be expressed as
$E_{\mathrm{LY}}=\frac{1}{8\pi}\int_{\Sigma}\left(H_{0}\,-\sqrt{H^{2}-({\rm tr}_{\Sigma}k)^{2}}\right)d\mu.$ (24)
They established positivity under the dominant energy condition (with $\Lambda=0$).
Theorem 3.2 ([LY03, LY06]).
Suppose $(\Omega,g,k)$ is a compact initial data set satisfying the dominant energy condition (with $\Lambda=0$) and has nonempty boundary $\Sigma=\partial\Omega$ of positive Gauss curvature and $H>|{\rm tr}_{\Sigma}k|$. Then the Liu–Yau energy $E_{\mathrm{LY}}$ of $\Sigma$ is non-negative and equals zero if and only if $(\Omega,g,k)$ can be embedded within Minkowski spacetime $\mathbb{R}^{3,1}$ and $\Sigma$ has only one component.
However, Ó Murchadha, Szabados, and Tod [MST04] constructed examples to show that there exist surfaces in Minkowski space satisfying these assumptions for which both the Brown–York and Liu–Yau masses are strictly positive.
To address this issue, Wang and Yau [WY09b, WY09a] introduced a new quasi-local energy based on isometric embeddings of 2-surfaces into Minkowski space. Their construction involves finding an embedding that minimizes a certain energy functional associated with the surface, and they proved existence and uniqueness results for such embeddings. This leads to a new expression of quasi-local mass for a large class of admissible surfaces. They showed that such mass is positive whenever the ambient spacetime satisfies the dominant energy condition, vanishes for surfaces in Minkowski space, and satisfies a rigidity property: if the mass is zero, the surface is isometric to one in flat space (see [WY09a]).
4. Supporting lemmas and proof of main results
In this section, we prove some supporting lemmas and Theorems 2.3 and 2.4. The first lemma provides a sufficient condition on the Gauss curvature of $\Sigma$ that ensures it can be isometrically embedded inside the cosmological horizon $r=1/\lambda$ in $\mathbb{R}^{3}$. The second and third lemmas address the positivity of the energy $E_{\lambda}$. Finally, we present the proofs of the main theorems by combining the lemmas.
4.1. Isometric embedding inside the cosmological horizon
The next lemma gives a sufficient (albeit not optimal) condition for $\Sigma$ to lie inside the cosmological horizon $r<1/\lambda$ once embedded into $\mathbb{R}^{3}$.
Lemma 4.1.
Let $\Sigma$ be a Riemannian two-sphere. For $\lambda>0$, suppose the Gauss curvature $K$ of $\Sigma$ is bounded below by
$K\geq\frac{3}{8}\pi^{2}\lambda^{2}.$ (25)
Then the isometric image of $\Sigma$ under the Weyl embedding into $\mathbb{R}^{3}$ is contained in the interior of a 3-ball with radius $1/\lambda$.
Proof.
In two dimensions, the Ricci tensor satisfies ${\rm Ric}=K\gamma$, where $\gamma$ is the metric on $\Sigma$. Using (25) and the Bonnet–Myers theorem, we obtain a bound on the diameter of $\Sigma$,
$\mathrm{diam}(\Sigma)\leq\sqrt{\frac{8}{3}}\,\frac{1}{\lambda}.$ (26)
Let $\Omega_{0}$ denote the compact and convex region enclosed by $\Sigma$ within $\mathbb{R}^{3}$ (determined by the Weyl embedding theorem). By convexity, the diameter of $\Omega_{0}$ [^2] To be precise, $\mathrm{diam}(\Omega_{0}):=\displaystyle\sup_{x,y\in\Omega_{0}}|x-y|$ while $\mathrm{diam}(\Sigma):=\displaystyle\sup_{x,y\in\Sigma}d_{\Sigma}(x,y)$. is strictly less than that of its boundary:
$\mathrm{diam}(\Omega_{0})<\mathrm{diam}(\Sigma).$ (27)
To see this, since $\Omega_{0}$ is compact and convex, its diameter is realized by a straight line in $\Omega_{0}$. That is, there are points $p,q\in\Omega_{0}$ such that $\text{diam}(\Omega_{0})=|p-q|$ which is the length of the straight line $c\colon[0,1]\to\Omega_{0}$ given by c(t)=tp+(1t)qc(t)=tp+(1-t)q. By convexity, $p$ and $q$ lie on the boundary $\Sigma$. Therefore
$|p-q|<d_{\Sigma}(p,q)\leq\text{diam}(\Sigma).$ (28)
The strict inequality in (28) follows since $K>0$. Indeed, seeking a contradiction, suppose $|p-q|=d_{\Sigma}(p,q)$. Then the image of $c$ must lie in $\Sigma$ (minimizing curves in Euclidean space are unique); hence $c$ is a pregeodesic in $\Sigma$. But, since $c$ is a geodesic in Euclidean space, it follows that the second fundamental form of $\Sigma$ within $\mathbb{R}^{3}$ vanishes when evaluated on $c^{\prime}$; this contradicts the fact that $\Sigma$ has positive Gauss curvature.
Jung’s theorem [Jun01] shows that $\Omega_{0}$ is enclosed in a ball of radius
$r=\sqrt{\frac{3}{8}}\,\mathrm{diam}(\Omega_{0}).$ (29)
Combining (26), (27), and (29), we conclude that $\Sigma$ is contained in a ball of radius $r<\frac{1}{\lambda}$, which completes the proof. ∎
Remark.
The curvature bound in equation (25) is not sharp. Indeed, consider a round sphere of radius $r=\frac{1}{\lambda+\varepsilon}$, for $\varepsilon>0$. Such a sphere clearly lies within the cosmological horizon. However, its Gauss curvature is $K=(\lambda+\varepsilon)^{2},$ which violates the bound in (25) whenever $\varepsilon$ is sufficiently small (note that $\tfrac{3}{8}\pi^{2}\approx 3.7)$. Indeed (25) is only a sufficient condition for our proof and the Jung’s theorem radius would be realized by a tetrahedron.
4.2. Positivity of energy
Our goal is to establish the positivity of the energy $E_{\lambda}$. As in Liu–Yau, we always assume $H>|\text{tr}_{\Sigma}k|$.
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with boundary $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ satisfying the dominant energy condition for $\Lambda=3\lambda^{2}\geq 0$.
Let $K_{\text{min}}=\min_{\Sigma}K$. We define
$\alpha_{0}=\sup_{C\in(0,1]}\min\left\{C\sqrt{K_{\text{min}}},\frac{1+\sqrt{1-C^{2}}}{C}\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right\},$ (30)
where $r_{\Sigma}$ is the area radius of $\Sigma$ and $E_{\rm LY}$ is the Liu–Yau energy of $\Sigma$ given by (24).
Note that $\alpha_{0}\geq 0$ since $E_{\rm LY}\geq 0$ (since $\mu\geq|J|_{g}$). Moreover, if $\mu>0$, the rigidity part of Theorem 3.2 implies that $E_{\rm LY}>0$ and hence $\alpha_{0}>0$; in this case, a little algebra shows that $\alpha_{0}$ is realized by a unique $C_{0}\in(0,1]$ and calculated below:
$\alpha_{0}=\begin{cases}\sqrt{K_{\min}}&\text{if }\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\\ \sqrt{\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right)}&\text{if }\sqrt{K_{\min}}\geq\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\end{cases}.$ (31)
The next lemma shows that for fixed geometric data, $E_{\lambda}\geq 0$ provided $\lambda$ is not too large.
Lemma 4.2.
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}\geq 0$ holds. If $\lambda\in[0,\alpha_{0}]$, where $\alpha_{0}$ is defined by (30), then the energy (15) is well defined and nonnegative:
$E_{\lambda}\geq 0.$ (32)
Moreover, if $0<\lambda<\alpha_{0}$, then $E_{\lambda}>0$, and if $E_{\alpha_{0}}=0$ with $\alpha_{0}>0$, then $\Sigma$ is isometric to a round sphere and $\sqrt{K_{\min}}\geq\frac{E_{\rm{LY}}}{r_{\Sigma}^{2}}$.
Proof.
For any $C\in(0,1]$, define
$\alpha_{C}=\min\left\{C\sqrt{K_{\text{min}}},\frac{1+\sqrt{1-C^{2}}}{C}\frac{E_{\rm LY}}{r_{\Sigma}^{2}}\right\}.$ (33)
Note that $\alpha_{0}=\sup_{C\in(0,1]}\alpha_{C}$. Moreover, if $\alpha_{0}>0$, then the map $(0,1]\mapsto(0,\alpha_{0}]$ given by $C\mapsto\alpha_{C}$ is surjective.
[…]
If $\lambda=0$, then $E_{\lambda=0}=E_{\rm LY}\geq 0$. On the other hand, if $\lambda\in(0,\alpha_{0}]$, choose $C\in(0,1]$ such that $\alpha_{C}=\lambda$. Then (34) shows that $E_{\lambda}$ is well defined and
$\begin{split}E_{\lambda}&=E_{\alpha_{C}}\\ &=\frac{1}{8\pi}\int_{\Sigma}\left(\sqrt{H_{0}^{2}-4\alpha_{C}^{2}}-\sqrt{H^{2}-({\rm tr}_{\Sigma}k)^{2}}\right)d\mu\\ &=\frac{1}{8\pi}\int_{\Sigma}\left(H_{0}-\sqrt{H^{2}-({\rm tr}_{\Sigma}k)^{2}}+\sqrt{H_{0}^{2}-4\alpha_{C}^{2}}-H_{0}\right)d\mu\\ &=E_{\rm LY}-\frac{1}{8\pi}\int_{\Sigma}\frac{4\alpha_{C}^{2}}{H_{0}}\frac{1}{1+\sqrt{1-4\alpha_{C}^{2}/H_{0}^{2}}}\\ &\geq E_{\rm LY}-\frac{|\Sigma|}{8\pi}\frac{2C\alpha_{C}}{1+\sqrt{1-C^{2}}}\\ &=E_{\rm LY}-\frac{C}{1+\sqrt{1-C^{2}}}\alpha_{C}r^{2}_{\Sigma}\\ &\geq 0.\end{split}$ (35)
In the second to last inequality, we used (34); in the last inequality, we used (33).
[…]
4.3. Proof of main results.
Combining the lemmas above, we can now prove Theorem 2.3.
Proof of Theorem 2.3.
By definition, we have $\alpha_{\lambda}\leq\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\rm min}}$ and so $\alpha_{\lambda}\leq\alpha_{0}$, where $\alpha_{0}$ is given by (31). Therefore, since $\lambda\leq\alpha_{\lambda}$ by hypothesis, Lemma 4.1 implies that we embed within the cosmological horizon and Lemma 4.2 implies $E_{\lambda}\geq 0$. The rigidity claim is obtained as in Lemma 4.2 except $C$ ranges from $0<C\leq\frac{2\sqrt{6}}{3\pi}$. ∎
Proof of Corollary 2.4.
We let $\lambda\geq 0$ be any constant such that $H>2\lambda$ and so
$E_{\rm LY}=\frac{1}{8\pi}\int_{\Sigma}(H_{0}-\sqrt{H^{2}-4\lambda^{2}})d\mu$ (36)
is defined. We aim to find a constant $\lambda_{\max}$ that does not depend on $\lambda$. For some constant $E\geq 0$ we define
$\alpha(E)=\begin{cases}\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}&\text{if }\sqrt{K_{\min}}<\frac{E}{r_{\Sigma}^{2}}\\ \min\left\{\sqrt{\frac{E}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E}{r_{\Sigma}^{2}}\right)},\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\min}}\right\}&\text{if }\sqrt{K_{\min}}\geq\frac{E}{r_{\Sigma}^{2}}\end{cases}.$ (37)
We observe that $\alpha_{\lambda}=\alpha(E_{\rm LY})$ is $\lambda$-dependent, while $\lambda_{\max}=\alpha(E_{\rm BY})$ is not. We claim that $\alpha(E_{\rm LY})\geq\alpha(E_{\rm BY})$. For these values of $\lambda$, note that $E_{\rm LY}\geq E_{\rm BY}$. If $\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}$, then the claim follows immediately from the definition (37). Suppose instead that
$\sqrt{K_{\min}}\geq\frac{E_{\rm LY}}{r_{\Sigma}^{2}}.$ (38)
As in the proof of Theorem 2.3, we define
$\alpha_{0}(E)=\sqrt{\frac{E}{r_{\Sigma}^{2}}\left(2\sqrt{K_{\min}}-\frac{E}{r_{\Sigma}^{2}}\right)}.$ (39)
Then a simple computation shows
$\begin{split}\alpha_{0}^{2}(E_{\rm LY})-\alpha_{0}^{2}(E_{\rm BY})&=\frac{2\sqrt{K_{\min}}(E_{\rm LY}-E_{\rm BY})}{r^{2}_{\Sigma}}-\frac{E_{\rm LY}^{2}-E_{\rm BY}^{2}}{r^{4}_{\Sigma}}\geq\frac{(E_{\rm LY}-E_{\rm BY})^{2}}{r_{\Sigma}^{4}}\geq 0,\end{split}$ (40)
where we used (38). This establishes the claim. Therefore, if we choose any $\lambda\in[0,\lambda_{\max}]$ where $\lambda_{\max}=\alpha(E_{\rm BY})$, we have $\lambda\leq\alpha_{\lambda}$ and so by Theorem 2.3, the energy $E_{\lambda}$ is nonnegative. Lastly, if $\lambda_{\max}>0$, then equation (37) implies $\alpha(E_{\rm BY})>0$ and so $E_{\rm BY}>0$. In this case, an argument as in (35) shows that $E_{\lambda}>0$ for all $0<\lambda\leq\lambda_{\rm max}$.
5. Examples
In this section we present some examples of solutions to the Einstein equations with positive cosmological constant that are foliated by umbilic slices.
5.1. Schwarzschild–de Sitter
In this section, we review the Schwarzschild–de Sitter spacetime, a $\Lambda$-vacuum solution to the Einstein equations. First, we show how to express the metric in flat expanding coordinates. Then we study the positivity of energy $E_{\lambda}$ along coordinate spheres. Lastly we determine $\alpha_{\lambda}$.
Fix a mass parameter $m>0$. Let $\Lambda>0$ denote the cosmological constant, and set as usual $\lambda=\sqrt{\tfrac{\Lambda}{3}}.$
Recall that the Schwarzschild–de Sitter metric in static coordinates is given by
$g=-\left(1-\frac{2m}{r}-\lambda^{2}r^{2}\right)\,dT^{2}+\frac{dr^{2}}{1-\frac{2m}{r}-\lambda^{2}r^{2}}+r^{2}\,d\Omega^{2},$ (41)
where $(S^{2},d\Omega^{2})$ denotes the unit round two-sphere.
If we perform the coordinate change
$T=t+\Phi(r),\qquad\Phi^{\prime}(r)=\frac{\lambda r}{\left(1-\frac{2m}{r}-\lambda^{2}r^{2}\right)\sqrt{1-\frac{2m}{r}}},$ (42)
then we obtain the metric in expanding coordinates:
$g\,=\,-\left(1-\frac{2m}{r}-\lambda^{2}r^{2}\right)\,dt^{2}-\frac{2\lambda r}{\sqrt{1-\tfrac{2m}{r}}}\,dr\,dt+\frac{dr^{2}}{1-\tfrac{2m}{r}}+r^{2}\,d\Omega^{2}.$ (43)
Recognize that when $m=0$ (43) is just the de Sitter metric (11) in flat-expanding coordinates.
Let $M_{t}$ denote a hypersurface of constant $t$ and $g$ its induced metric. If we denote the second fundamental form of $M_{t}$ by $k$, following our convention (1), a straightforward computation shows umbilicity:
$k=\lambda g.$ (44)
Consider a large coordinate sphere $\Sigma_{R}=\{r=R\}$. Let $H$ denote its mean curvature within $M_{t}$ and $H_{0}$ its mean curvature within Euclidean space via the Weyl embedding theorem. Assume $R$ is chosen such that
$H=\frac{2}{R}\sqrt{1-\frac{2m}{R}}>2\lambda,\qquad H_{0}=\frac{2}{R}>2\lambda.$ (45)
We have
$\begin{split}E_{\lambda}(\Sigma_{R})&=\frac{R^{2}}{2}\left(\sqrt{\frac{4}{R^{2}}-4\lambda^{2}}\,-\,\sqrt{\frac{4}{R^{2}}\left(1-\frac{2m}{R}\right)-4\lambda^{2}}\right)\\ &>E_{\rm BY}(\Sigma_{R})\\ &>0.\end{split}$ (46)
Therefore, the energy is positive for all the values of $\lambda$ for which it is well defined.
Note now that since $K=\frac{1}{R^{2}}$,
$\begin{split}\frac{E_{\rm LY}(\Sigma_{R})}{R^{2}}&=\frac{1}{R}\left(1\,-\,\sqrt{1-\frac{2m}{R}-\lambda^{2}R^{2}}\right)<\sqrt{K_{\min}}.\end{split}$ (47)
Hence,
$\alpha_{\lambda}=\min\left\{\sqrt{\lambda^{2}+\frac{2m}{R^{3}}},\frac{2\sqrt{6}}{3\pi R}\right\}.$ (48)
Note that the second expression in the definition of $\alpha_{\lambda}$ was chosen to guarantee, via Lemma 4.1, that the embedding of $\Sigma_{R}$ lies inside the cosmological horizon $r=\tfrac{1}{\lambda}$. This requirement, however, is not sharp: in the present setting it suffices to have $R\leq\frac{1}{\lambda}$, which is ensured by (45). Thus only the first expression in 48 needs to be considered. Clearly,
$\lambda\leq\sqrt{\lambda^{2}+\frac{2m}{R^{3}}.}$ (49)
Therefore $E_{\lambda}>0$ for all $\lambda$ whenever (45) holds.
The results obtained in this section for the exact Schwarzschild de Sitter case, pose a physically meaningful stability question. That is, whether a similar statement would hold for perturbations of this example. By the above computations, this would appear to be related to an associated stability analysis for the Weyl embedding theorem, a problem which could be of interest on its own right.
5.2. Other $\Lambda$-vacuum examples
In the previous section, we applied Corollary 2.4 to the Schwarzschild–de Sitter spacetime, which is one example of a $\Lambda$-vacuum spacetime. In this section we show that there are many $\Lambda$-vacuum spacetimes where Corollary 2.4 applies.
Let $V$ be a closed Riemannian manifold with positive scalar curvature. Fix $p\in V$ and set $M:=V\setminus\{p\}$. The Green’s function at $p$ for the conformal Laplacian on $V$ exists and is strictly positive. Consequently there is a scalar flat metric $g$ on $M$ which is asymptotically flat with $p$ representing the point at infinity, see e.g., [Sch89, LP87]. Since $R_{g}=0$, the initial data set $(M,g,k=\lambda g,\lambda)$ satisfies the $\Lambda$-vacuum Einstein constraint equations with $\Lambda=3\lambda^{2}$. In the asymptotically flat coordinates $x,y,z$, let $\Sigma_{r}$ be the sphere given by $r^{2}=x^{2}+y^{2}+z^{2}$. For $r$ large enough, $\Sigma_{r}$ has positive Gauss and mean curvatures and bounds a compact region $\Omega_{r}$. Let $\lambda_{\rm max}$ be given by Corollary 2.4. Then for all $\lambda\in[0,\lambda_{\rm max}]$, we can choose $r$ large enough so that $E_{\lambda}$ of $(\Omega_{r},g,k=\lambda g,\lambda)$ is well defined and nonnegative.
6. Discussion and conclusion
In this work, we addressed the problem of defining energy within the framework of an initial data set admitting a positive cosmological constant. In particular, the presence of a cosmological horizon in the background de Sitter spacetime constitutes an obstruction to a global notion of energy, thereby motivating us to use quasi-local definitions instead. Inspired from the Liu–Yau construction in the asymptotically flat setting, we adapted their definition to the flat expanding patch of de Sitter spacetime:
$E_{\lambda}(\Sigma)=\frac{1}{8\pi}\int_{\Sigma}\left(\sqrt{H_{0}^{2}-4\lambda^{2}}-\sqrt{H^{2}-(\operatorname{tr}_{\Sigma}K)^{2}}\right)d\mu,$
where $\lambda=\sqrt{\Lambda/3}$ denotes the expansion parameter associated with the cosmological constant $\Lambda>0$. We established sufficient conditions guaranteeing that the isometric embedding of $\Sigma$ lies entirely within the cosmological horizon $r<1/\lambda$, and proved positivity of $E_{\lambda}$ for a broad class of initial data sets, including the perfectly umbilic case, $k=\lambda g$, for all sufficiently small $\lambda$.
Several open questions naturally arise from this framework. A first one concerns rigidity: by analogy with the classical Positive Energy Theorem, one expects that the vanishing of the quasi-local energy should characterize data isometric to a domain in the flat expanding de Sitter spacetime. Formally, one may conjecture that
[…]
So far, we have obtained only a partial rigidity result: in the rigid case, the boundary $\Sigma=\partial\Omega$ must be round, and this occurs precisely at the critical value $\lambda=\alpha_{\lambda}$. Another open problem concerns the dependence of the quasi-local energy on the parameter $\lambda$. As discussed in Theorem 2.3, our result can be interpreted as identifying a range of admissible values of $\lambda$ for a fixed initial data set $(\Omega,g,k)$. It would be interesting to further investigate how the parameter $\lambda$ interacts with the intrinsic and extrinsic geometry of $\Sigma$. A further promising direction is the development of an optimal embedding formulation. In the asymptotically flat case, Wang and Yau [WY09b, WY09a] introduced a variational quasi-local energy obtained by minimizing over admissible isometric embeddings into Minkowski space. Constructing an analogous variational framework in the de Sitter setting would provide a natural and geometrically consistent extension of our approach.
Acknowledgments
The research of EL was supported by Carlsberg Foundation CF21-0680 and, together with AP, by Danmarks Grundforskningsfond CPH-GEOTOP-DNRF151. RA would like to thank the Alexander von Humboldt foundation for partial financial support. This project took shape while EL and AP were participants in the program “New Frontiers in Curvature: Flows, General Relativity, Minimal Submanifolds, and Symmetry” at the Simons Laufer Mathematical Sciences Institute (SLMath), supported by Grant No. DMS-1928930 during the Fall 2024 semester. We are grateful for the many enlightening discussions with experts during the program, and in particular to Dan Lee for his insightful conversations. EL and AP also acknowledge the support of the Simons Center for Geometry and Physics, where parts of this work were carried out during the program “Geometry and Convergence in Mathematical General Relativity”. Lastly, we are deeply grateful to Pengzi Miao for valuable discussions and suggestions, especially regarding Lemma 4.1.